Displaying reports 81-100 of 1079.Go to page 1 2 3 4 5 6 7 8 9 10 End
Reports until 12:25, Tuesday 08 April 2025
Atoms General (RbQ)
Daniel Grass - posted 12:25, Tuesday 08 April 2025 (12228)
RbQ Ideal Vacuum Cubes Joined, Plates Added, and Mounted on Machined Breadboard

[Jeff, Ian, Daniel]

Ian and I machined out 1/4-20 holes in the Thorlabs MB1824 aluminum breadboard into 0.266" diameter through holes for 1/4 screws so that we can put screws into the blind holes on the Ideal Vacuum base plates. We used the drill press where we could and a drill otherwise for holes too close to the center of the breadboard. I accidentally flipped the orientation of the holes, so I am currently building the mirror image of my SolidWorks design. I don't think this will be an issue. If it is, we can machine more holes into the MB1824 breadboard.

Jeff and I cleaned with isopropanol wipes and attached the two ideal vacuum "cubes" following these steps. We put the "high pressure side" away from the MOT because I figured it would be easiest to screw in parts there. The cubes were flush together, but the cube was oriented a bit clockwise when looking down from the cube to the prism.

We then attached the sides to the cube/prism assembly. For the 6"x6" plate to 2" diameter window, we cleaned and dropped in one viton gasket but couldn't fit a second gasket between the window (which we cleaned with puffs of clean air) and the "screw", partially because the mirror is too thick, partially because the viton is too wide when it's on the "screw". I tightened the "screw" finger tight. This can't be good for the window, but the edge of the window can't be used anyway. The "screw" sticks out approximately 0.75-0.76" for all 3 we assembled.

For tightening the plates to the cube/prism assembly, I did a few turns by finger to get the screws in, then used a screw driver until I encountered some resistance, then did a bit more on the screwdriver, before finishing with the torque wrench at 50 in lbs (5.7 Nm) and then 70 in lbs (7.9 Nm) twice, following these instructions. A torque wrench that can be set lower (like this) and corresponding hex bit socket set, if needed, (like this) could be useful. I used a lot of isopropanol wipes since the cleanliness of the parts is suspect. Hopefully everything is sufficiently clean.

Lastly, Jeff and I mounted the prism/cube assembly to the breadboard with some 3/4" long 1/4-20 screws and washers.

Next, I am going to attach parts to the CF flanges on the ideal vacuum plates. I am not sure which viton gasket I should use. (1 or 2)

Images attached to this report
Vacuum Systems General (GQuEST)
Daniel Grass - posted 23:04, Thursday 03 April 2025 (12225)
Blank 4.5 in CF Flange and 2.75 in CF Flanges attached to cleaned Zero Length Reducer Flanges on Demonstrator IFO Vacuum Cube

I cleaned the 10" to 4.5" CF Zero Length Reducer Flange and the two 10" to 2.75" CF Zero Length Reducer Flanges on the Demonstrator IFO Vacuum Cube from duct tape and its residue with wipes, acetone, and a bit of isopropanol. The 10" to 4.5" CF Zero Length Reducer Flange has some grime that I partially removed. I tightened the bolts on the 4.5" CF Flange to 27 Nm which seemed to give a steel to steel contact such that the copper gasket wasn't visible. I tightened the bolts on the 2.75" CF Flanges to 18.6 - 19 Nm which gave a steel to steel contact. Next steps are to install the 10" to 6" Zero Length Reducer Flange, the 6" 6 way cross and all of the vacuum parts off of it.

GQuEST Output Filter Cavities
Torrey Cullen - posted 16:00, Thursday 03 April 2025 (12224)
Stable Piezo Lock in OFC3

I set up the piezo control for the third filter cavity. This is to allow a way to scan cavity 3 while keeping OFC2 laser locked on 1550. The scan works and after tuning the controller a little it provide one of the best piezo locks we've seen in these cavities. Note that this new cavity is with the new mirror mounts and the nac2125 from CTS piezo. More analysis is needed but this piezo seems promising. For now I want to align the OFC3 1550 path, which I now should have all the tools for. Will come back to this new piezo analysis later.

GQuEST Output Filter Cavities
Torrey Cullen - posted 14:19, Thursday 03 April 2025 (12223)
OFC2 OFC3 inter cavity mode matching

Continuation of this work.  Yesterday, Jeff and I set up a way to have constant light exiting OFC2 to align the 1550 path for OFC3. I placed lenses according to the mode matching solution in the previous post and profiled the light coming out. Plugging in these numbers yields less than 1% mode mismatch for cavity 3. I will attempt to align 1550 light into the cavity now.

Images attached to this report
Electronics General
Jeffrey Wack - posted 20:06, Wednesday 02 April 2025 - last comment - 15:42, Monday 07 April 2025(12221)
Power supply for Thorlabs laser

I tried to power the ThorLabs ULN15TK with a UX4 EXCELSYS power supply. It seems as though we have the XgF/Xg8 powerMods installed, each with a max current draw of 3A, so I wired two modules together in parallel to surpass the 4A asked for by the laser.

Plugging in the laser results in a green power light and no status light, but the emission light does not turn on after hitting enable.

We should perhaps buy another DS12 to power the ULN15TK, but I remain confused as to why the other power supply did not work. It is providing the correct voltage, but perhaps wiring two modules in parallel does not increase the current supply as I expect?

The state of the lasers is that both lasers and the amplifier are off. The signs remain on.

Comments related to this report
Daniel Grass - 10:46, Thursday 03 April 2025 (12222)

Laser seeder and amplifier are back on

Jeffrey Wack - 15:42, Monday 07 April 2025 (12227)

The new DS12 arrived in the mail. After using it to power the ThorLabs laser, we found that the 900Hz noise is present, but at a much smaller magnitude compared to the old supply.

Images attached to this comment
Electronics General (GQuEST)
Alex Ramirez - posted 18:53, Wednesday 02 April 2025 (12220)
PCBs for Dewar Thermometry

I have assembled and soldered the Thermometry PCBs to be mounted into the Dewar for initial pump-down tests. These PCBs were soldered in the vent hood in the chem room with lead-based solder and acuu glass flux that is vacuum compatible.

Upon finishing the boards, they were then soaked in isopropyl alcohol and sonically cleaned for 5 minutes (as directed on the accu glass website). 

Please see the last photos for side-by-side comparison of post soldering and post-sonicating.
 

Images attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 17:16, Wednesday 02 April 2025 (12217)
PDH setup for aux 1550 path to filter cavity 2

[Jeff, Torrey]

We setup a fiber EOM and a REFL PD to set up a PDH lock on the auxillary 1550 path to the second filter cavity. This will give consistent 1550 light leaving filter cavity 2 in order align and mode match into filter cavity 3.

First, we added a mirror to direct the reflected 1550 beam to a Newport 1811 photodiode. This mirror is cramped against the imput mirror and is very close to clipping the input beam. We used a lens and an ND filter to focus onto the PD and prevent saturation. The REFL dip is not very low, suggesting the cavity is not critically coupled or we have alignment issues. Additionally, the old 900ish Hz oscillations are seen on this photodiode signal. Even though we are not using the Thorlabs laser, this is expected because the whole drawer is sharing ground with the Thorlabs laser. Hopefully this can be resolved by swaping out the Thorlabs power supply.

Next we installed a fiber EOM on the OFC 2 1550 test path. The RF drive required a SMP connector, which was borrowed from the LIGO lab from Todd Etzel. 

There was initially a large optical loss through the EOM, this was addressed by adding a half waveplate at the fiber input of the test path on the power distribution sled and rotating the polarization to maximize transmitted power.

After tuning up alignment a bit, we achieved a lock of OFC 2 on 1550, actuating on the Teraxion laser.

We unplugged the Thorlabs laser so it is completely off, and this removed to 900Hz noise and improved the lock.

Electronics General
Jeffrey Wack - posted 16:20, Wednesday 02 April 2025 (12216)
AOM driver power supply consolidation

Both AOM driver chassis are now being powered by the same 'excelsys UX4' supply. It provides +24V and GND to the RF amplifiers as depicted in 12179.

After the change I checked that all 3 AOMs are still correctly deflecting a beam, and a twice shifted beam ends up at the fiber collimators.

Images attached to this report
Vacuum Systems General (GQuEST)
Daniel Grass - posted 12:18, Wednesday 02 April 2025 (12215)
Turbo and Ion Pump Controllers fully Connected Except to Wall Outlet

I connected the Agilent TwisTorr 74 Rack Controller to the TwisTorr 74 Turbo Pump, its cooling fan, and the FRG 702 pressure gauge. I also connected the ion pump controller to the ion pump. In case there is an accidental power on, I am waiting to plug them into a wall outlet until I am ready for them to be on.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 12:17, Wednesday 02 April 2025 (12214)
2.75 in 1550 nm AR Coated Viewport installed on Laser Filter Cavity Input Vacuum Cube

I installed a 2.75" 1550 nm AR Coated Viewport on the Laser Filter Cavity (LFC) Output Vacuum Cube. I tightened the bolts to 10.7 Nm. 9.6 Nm was not enough to get a full steel to steel contact. I checked the other viewport I installed, and I could see a bit of a gap. I therefore tightened those bolts to 10.7 Nm as well and the gap reduced in size. Other than the hose from the turbo pump to the scroll pump, the LFC should be completely sealed and ready to be pumped out.

Images attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 12:03, Wednesday 02 April 2025 (12213)
Intercavity Mode Matching (OFC2 -> OFC3)

With the 775 OFC3 cavity lock successful, it is time to work on 1550. This requires mode matching the cavity. The light is diverging when it exits the cavity, with a waist of 581 um, roughly .6 m before the exit of the cavity (on M1 of OFC2). I adjusted the inter cavity mirrors position to maximize the amount of the length the light can travel as this is better for mode matching in this case. The two intercavity mirrors are at 15 and 18 inches from the cavity exit. The OFC3 cavity input mirror is at 29 inches. We want a waist here of 577.7 um (according to my finesse calculations of the cavity). JAMMT spits out an f=2m and f=1m lens solution based on the lens we have available. I've assembled these lens, moved the mirrors back into their positions, roughly placed them for now. See attached image for the calculation.

I am going to attempt to align into the next cavity using OFC2 1550 light. This poses a challenge because you need to make the two wavelengths coresonant and then lock on 775. Because the bandwidth of the 1550 cavity is so low and the lock quality is poor (and there seems to be 1550 -> 775 drift via the aom) it makes the test light being used very unstable. Will attempt this after lunch.

Images attached to this report
Vacuum Systems General
Daniel Grass - posted 22:08, Tuesday 01 April 2025 (12212)
5/16 in 12-point star Torque Adapter Wrench Sanded Down for Use on the Agilent TwisTorr 74 Turbo Pump

I used a belt sander to reduce the diamter around the 5/16" 12-point star Torque Adapter Wrench for Use on the Agilent TwisTorr 74 Turbo Pump since the turbo pump comes quite close to the screws, preventing the wrench from easily fitting around them.

The screw diamter is ~0.462-0.468". I ground down some of the wrench so that its radius was smaller than screw's. I confirmed the fit of the wrench around the screw that doesn't hit the pump.

Equipment Loans (GQuEST)
Daniel Grass - posted 22:01, Tuesday 01 April 2025 (12211)
Discussion on 4D Technology NanoCam HD Equipment Loan

I spoke with GariLynn Billingsley and Rodica Martin today about using the 4D Technology NanoCam HD to image the GQuEST End Mirrors with the surface correction mount. GariLynn understandably did not want the expensive profiler to be in Bridge instead of their custody, but said that we could continuously use it over the summer somewhere in Downs-Lauritsen with the exception of a few days when they need it. They think they can mount it on some optics table, although we hopefully will be looking for ~1 nm precision so the vibration isolation will be less important.

I am trying to figure out if there is a good way to mount the profiler so that it looks parallel to the ground instead of down.

Lab Infrastructure General
Ian MacMillan - posted 16:24, Tuesday 01 April 2025 (12210)
Key Box

I hung the box with the keys on the wall outside the GQuEST Lab under the light-up laser sign. See Key_box_location.png for location. The code for the box can be found in the lab secrets wiki. The information on the keys inside the box can be found on the key box wiki page. These are backup keys most of the things that have keys in this box are left unlocked. If you have a key for the box, add it to the paper in the box and to the key box wiki.

Images attached to this report
Lab Infrastructure General
Ian MacMillan - posted 15:35, Tuesday 01 April 2025 (12209)
Lista Cabinets Caster Platforms

[Ian, Alex, Torrey]

We assembled the platforms with casters on which the Lista cabinets will be placed. As seen in Lista_Wheels.jpeg, there are two fixed wheels and two swiveling wheels on each platform. The next step is to get facilities to lift up the Lista cabinets and place them on the platforms.

Images attached to this report
SNSPD General (GQuEST)
Daniel Grass - posted 17:12, Thursday 13 February 2025 - last comment - 15:04, Saturday 05 April 2025(12137)
Aluminum KF50 Centering Rings Machined to hold PCB in SNSPD Dewer

[Alex, Daniel]

I got two aluminum KF50 Centering Rings from Nick Hutzler's group and machined them in the lathe to remove one of their lips so that they can hold a PCB that interfaces with the inside and outside of the Dewer. I clamped them on the outside lip with a 6 jaw chuck. I used a moderate amount of clamping force (~90 degrees of rotation with the chuck key) to hold them. I slowly increased the spindle speed to ensure the rings wouldn't fly off; I used a final spindle speed of 1700 rpm, which is pretty standard for aluminum and a carbide tool. I kept the x-axis of the lathe at around 1.93" and slowly moved the tool in the z-axis to remove the lip. I needed to take off 0.080" of material. For the last few thousands of an inch, I moved the tool in the x-axis instead of the z-axis. I then used a deburring tool by hand to remove any burrs from machining. I could not see any markings from the clamping jaws.

Tomorrow, I will clean and start to bake out these centering rings.

Comments related to this report
Daniel Grass - 12:32, Friday 14 February 2025 (12138)

We cleaned the centering rings with the normal aluminum procedure of successive 3 minute baths and scrubs of 1:30 Simple Green:DI water, DI water, and isopropanol. We put it in the vacuum oven to bake out at 120° for 48 hours since we don't know the alloy of aluminum. 

Daniel Grass - 21:39, Tuesday 18 February 2025 (12141)

I turned off the vacuum pump and closed the valve to the vacuum pump around noon today. According to the vacuum oven, the bakeout lasted 49-50 hours. We can store these parts under vacuum unitl we need them or the vacuum oven.

Daniel Grass - 17:32, Tuesday 04 March 2025 (12161)

I machined 3 more aluminum KF50 centering rings exactly like I did the first 2.

Daniel Grass - 18:19, Wednesday 02 April 2025 (12219)

We moved the two cleaned centering rings out of the oven and put them in UHV foil and an antistatic bag. We repeated the cleaning process for the three dirty rings and are baking them out.

Daniel Grass - 15:04, Saturday 05 April 2025 (12226)

I turned off the vacuum pump and closed the valve to the vacuum pump

Lab Infrastructure General (GQuEST)
Daniel Grass - posted 19:18, Monday 11 November 2024 - last comment - 18:17, Wednesday 02 April 2025(11976)
Guardrails made and installed onto shelf above optics tables

[Ian, Sander, Daniel]

I designed some guardrails that go 3" above the M6 tapped holes of the shelf above the optics tables in B111B. I used 1/8" thick, 1/2" wide aluminum bar. There were 3 piece types (all holes centered along the 1/2" width):

Vertical bar: 4" long; 0.24" (C drill) diameter hole 0.5" from bottom; 0.27" (H drill) diameter hole 3.5" from bottom

Long horizontal bar: 36" long; 0.27" diameter holes at 0.78", 1.5", and 3 more holes with a regular 250 mm spacing beyond 1.5" to match the shelf hole pattern.

Short horizontal connection bar: 19" long; 0.27" diameter holes 0.375" from side and 2 more holes with 9.125" spacing

This short horizontal connection bar is designed to be replaced by a patch panel if desired.

Ian helped me drill some holes and Sander helped me install the connection bars.

We were not able to install 1 long and 1 short horizontal bar because the seeder breadboard overhangs the shelf by ~1". 

 

See attached photos.

Images attached to this report
Comments related to this report
Daniel Grass - 18:17, Wednesday 02 April 2025 (12218)

[Jeff, Daniel]
 

I drilled a few more parts and installed the last pieces of the guardrail since we moved the breadboard.

Displaying reports 81-100 of 1079.Go to page 1 2 3 4 5 6 7 8 9 10 End